### v1

# OmniDRCA: Parallel Speech-Text Foundation Model via Dual-Resolution Speech Representations and Contrastive Alignment

Abstract
Recent studies on end-to-end speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens. Existing approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis. (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation. This paper presents OmniDRCA, a parallel speech-text foundation model based on joint autoregressive modeling, featuring dual-resolution speech representations and contrastive cross-modal alignment. Our approach processes speech and text representations in parallel while enhancing audio comprehension through contrastive alignment. Experimental results on Spoken Question Answering benchmarks demonstrate that OmniDRCA establishes new state-of-the-art (SOTA) performance among parallel joint speech-text modeling based foundation models, and achieves competitive performance compared to interleaved models. Additionally, we explore the potential of extending the framework to full-duplex conversational scenarios. [^1] Our implementation will be released at https://github.com/lxchtan/OmniDRCA.
1 Introduction
Recent advancements in Large Language Models (LLMs) (OpenAI, 2023; Yang et al., 2024) have demonstrated remarkable performance across various NLP tasks, sparking growing interest in extending their capabilities to multimodal domains encompassing vision and audio. Developments in spoken dialogue systems are critical to human-computer interaction, as natural human communication inherently relies on verbal exchanges. The release of GPT-4o (OpenAI, 2024b) is a significant milestone in this field, showing the potential of LLM-based spoken dialogue systems for seamless and natural interactions with users.
LLM-based spoken dialogue systems can be generally categorized into cascaded systems and end-to-end (E2E) systems, with the distinction lying in whether the backbone LLM can directly comprehend speech representations and generate speech outputs. Early cascaded systems integrate separately trained Automatic Speech Recognition (ASR), LLM, and Text-to-Speech (TTS) modules. Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., tone, emotion), and significant latency. Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (Chu et al., 2023, 2024). E2E systems have emerged to further eliminate ASR and TTS modules, and establish direct connections between speech representations and LLMs. However, training LLMs to generate highly intelligent speech outputs remains challenging. E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
Recent research efforts on E2E models to address the above challenges have focused on two primary directions: Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models. Text-Driven Speech Models refer to systems in which LLMs ingest speech representations as input, produce textual responses, and simultaneously utilize the LLM’s latent representations to synthesize audio. In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and autoregressively generating both text tokens and speech tokens. The key distinction lies in whether the generated speech tokens are fed back autoregressively into the LLM during the generation process. Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024). Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM, ensuring temporal alignment. While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations. Text-Driven Speech Models prevent LLMs from perceiving generated speech tokens during inference, thereby limiting dynamic adjustment of text generation. Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge. Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials; hence, in this work, we focus on enhancing joint speech-text modeling.
In this work, we propose OmniDRCA, a novel parallel speech-text foundation model. Our architecture preserves end-to-end advantages while significantly improving speech-text semantic alignment through two key innovations: 1) Dual-Resolution Speech Representations (DRSR): DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens. For the LLM’s speech comprehension, we introduce a grouping mechanism, mapping 25Hz discrete audio tokens to 5Hz speech representations that align much better with textual granularity, thereby leveraging LLM’s capabilities more effectively and improving speech understanding. Combined speech-text embeddings form the assistant’s autoregressive input. The multimodal LLM (MLLM) predicts text tokens. For speech generation, latent representations are ungrouped and autoregressively processed by the speech refinement model (SRM) for individual speech token prediction. During inference, due to input-output resolution discrepancies, each MLLM inference triggers $k$ parallel SRM executions, where $k$ is the grouping factor. 2) Contrastive Cross-modal Alignment (CCA): Similarity between grouped speech representations and text representations is computed to optimize their mutual distances. While grouping aligns speech-text granularity, CCA further enhances semantic consistency between modalities before they enter the LLM and further enhances speech comprehension. Considering inference efficiency for varying output requirements in practical applications, we also design system prompts to control the output modes, including text-only output, parallel speech-text output, and parallel speech-text output after chain-of-modality (CoM) (Zhang et al., 2023a). Experimental results demonstrate the efficacy of CoM in improving performance. For reproducibility, we will release all source code and model checkpoints.
Our contributions are three-fold: 1) We propose OmniDRCA, a novel parallel speech-text foundation model that enhances speech-to-speech generation based on discrete speech tokens and LLMs, through dual-resolution speech representations and contrastive cross-modal alignment. 2) Comprehensive experiments demonstrate OmniDRCA’s superiority over the state-of-the-art (SOTA) parallel model LUCY (Gao et al., 2025) and competitive performance against SOTA interleaved models across multiple spoken question answering benchmarks. 3) We further explore the potential of OmniDRCA for full-duplex voice conversation. By leveraging time-division multiplexing, OmniDRCA-Duplex supports real-time responses to user interruptions with minimal performance degradation.
2 Related Work
Speech Tokenization. Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations. Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation. In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as in CosyVoice (Du et al., 2024a, b), and prioritized for linguistic content (Hsu et al., 2021). While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023). In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to its competitive semantic capabilities and support for streamlined streaming speech generation.
E2E Speech Foundation Models. Existing E2E speech foundation models can be categorized into Text-Driven Speech Models and Joint Speech-Text Models. Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech. The contemporaneous Text-Driven model Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production. Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
Joint Speech-Text Models explore two different architectures, Interleaved and Parallel. Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024) conduct parallel decoding. The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations. Our approach differs from Moshi with two key distinctions: (1) the $k$ tokens correspond to grouped semantic speech tokens rather than disparate acoustic tokens, and (2) our speech refinement module explicitly ingests full-sequence contextual information to enhance semantic coherence. While SLAM-Omni (Chen et al., 2024) proposes grouped semantic token prediction for speech generation, our approach designs an ungrouping mechanism and incorporates a Speech Refined Model for autoregressive generation of individual speech tokens. Another contemporaneous parallel model Kimi-Audio (KimiTeam et al., 2025) is closest to our approach. To mitigate limitations of discrete tokens in audio understanding, they introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper features, preserving both semantic and acoustic information. In contrast, our approach enhances discrete token comprehension through contrastive loss and grouping strategies, thereby improving LLM’s understanding of speech tokens while maintaining representational efficiency.
Contrastive Cross-Modal Alignment. CLIP (Radford et al., 2021) aligns image and text representations through contrastive pretraining. VILA-U (Wu et al., 2024) integrates contrastive loss and reconstruction loss during vision token pretraining to enhance feature alignment. In speech research, Ye et al. (2022) leverages multi-class N-pair contrastive loss (Sohn, 2016) for speech translation. Zheng et al. (2024) leverages InfoNCE loss (He et al., 2020) for spoken language understanding. We incorporate contrastive loss to strengthen speech-text semantic alignment. Crucially, we adopt gradient stop (Chen and He, 2021) on text embeddings during contrastive loss computation – this modification preserves speech representation learning while preventing text feature distortion.
3 Methodology
Figure 1 illustrates the architecture of OmniDRCA. The system consists of two main components: (1) a Multimodal Large Language Model (MLLM) designed to process interleaved speech-text inputs and achieve cross-modal alignment, and (2) a Speech Refined Model (SRM) specialized in high-fidelity speech synthesis. The system operates through multimodal input encoding, joint representation learning with contrastive alignment, and coordinated speech-text output generation. During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce multimodal context embeddings, and decoded into synchronized speech-text responses through SRM and the language model head.
3.1 Speech Tokenization and Detokenization
Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ HuBERT (Hsu et al., 2021) as the speech tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$, where $T$ is the speech token sequence length. For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency.
Figure 1: Overview of OmniDRCA. User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text token prediction. Hidden representations are ungrouped and processed by the SRM through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor. The generated speech tokens are then converted to speech waveform by the speech detokenizer. We omit speech tokenizer and detokenizer for simplicity.
3.2 Multimodal Large Language Model (MLLM)
Built upon text-LLMs, the MLLM learns joint speech-text modeling to process speech or text inputs while producing parallel speech and text outputs.
Parallel Joint Speech-Text Model. Leveraging the autoregressive generation capability of MLLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the MLLM at each timestep. Their embeddings are added as model input, forming a parallel speech-text architecture. Formally, the combined input embedding $c_{t}$ at timestep $t$ is computed as:
$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. The autoregressive generation process is as follows:
$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
where $x$ is the input sequence and $y_{t}=(s_{t},t_{t})$ denotes the joint speech-text output at timestep $t$. This unified formulation enables seamless integration of speech and text generation within a single autoregressive framework. To preserve the intrinsic linguistic understanding and generation capabilities of pretrained text-LLMs while enabling cross-modal interactions, three key methodological components are introduced for intermodal coordination.
Inner Text Streaming. Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold. We focus on performing modality alignment exclusively at the assistant end. This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
Speech Token Grouping. A crucial parameter for discrete tokens is sampling rate, which determines the input/output sequence length. GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz. Hence, our work adopts 25Hz sampling rate. To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024), a grouping mechanism is designed:
$\mathbf{g}_{i}=\text{Linear}\left(\mathop{\|}_{j=ik}^{(i+1)k-1}\mathbf{s}_{j}\right)\in\mathbb{R}^{d_{\text{text}}}$ (3)
[…]
Cross-modal Contrastive Alignment.
To enable cross-modal alignment, a contrastive learning objective aligns grouped speech representations with their textual counterparts by maximizing their cosine similarity and computing a contrastive loss. This joint optimization ensures coordinated positioning of both modalities within the shared representation space.
$\mathcal{L}_{\text{cont}}=-\cos(E_{\text{text}}(\mathbf{t}),E_{\text{speech}}(\mathbf{g})),$ (4)
where $E_{\text{text}}$ and $E_{\text{speech}}$ are the mean of embeddings of text tokens and grouped speech tokens between sample pairs, respectively. Different from early applications of contrastive alignment between speech and text (Zheng et al., 2024), our approach eliminates negative sample computation and prevents mode collapse through gradient stop on the text branch.
3.3 Speech Refined Model (SRM)
SRM is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the MLLM as conditional input and incorporates contextual speech information to autoregressively generate speech tokens. While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details. To address this limitation, an ungrouping process is performed as follows: the MLLM’s final hidden state is mapped to group-sized embeddings via linear projection
$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[MLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (5)
and time splitting
$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}]\quad\text{where}\quad\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k},$ (6)
the resulting $\mathbf{H}$ serves as conditional input for SRM that autoregressively generates speech tokens. Following our practice on MLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated. SRM is trained to maximize the conditional probability:
$\mathcal{L}_{\text{SRM}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}),$ (7)
where $s_{i}$ represents the $i$-th speech token. By optimizing this objective, SRM learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from MLLM. This design enables SRM to effectively leverage the semantic and acoustic information encoded in the hidden representations of MLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
The end-to-end training objective integrates the three losses through multi-task learning paradigm:
$\mathcal{L}_{\text{total}}=\lambda\mathcal{L}_{\text{LM}}+\mu\mathcal{L}_{\text{cont}}+\gamma\mathcal{L}_{\text{SRM}},$ (8)
$\mathcal{L}_{\text{LM}}=-\sum_{i=1}^{T}\log P(t_{i}|t_{<i},\mathbf{g}),$ (9)
where $\mathcal{L}_{\text{LM}}$ is the autoregressive loss over MLLM, $\lambda$, $\mu$, and $\gamma$ are hyperparameters.
3.4 Multimodal Interaction Patterns
Table 1: Multimodal Interaction Patterns.
Pattern Name Abbr. Modality Flow
Speech-to-Multimodal S2M Speech $\rightarrow$ Joint speech-text response
Speech-to-Text S2T Speech $\rightarrow$ Text-only response
Text-to-Multimodal T2M Text $\rightarrow$ Joint speech-text response
Text-to-Text T2T Text $\rightarrow$ Text-only response
Speech-Text Chain STC Speech $\rightarrow$ Text transcription $\rightarrow$ Text response $\rightarrow$ Multimodal response
Speech-Assisted Chain SAC Speech $\rightarrow$ Text response (agent perspective) $\rightarrow$ Multimodal response
Speech-User Chain SUC Speech $\rightarrow$ Text transcription (user perspective) $\rightarrow$ Multimodal response
Practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Furthermore, Zhang et al. (2023a) found that the chain-of-modality (CoM) methodology can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in the Appendix. For training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
4 Experiments
Table 2: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks. Baseline results are from Yao et al. (2024) except LUCY (S2) and GLM-4-Voice-Small from Gao et al. (2025) and Zeng et al. (2025) respectively. Metrics are benchmark-specific (higher is better). The Avg. score averages four benchmarks (Gray: AlpacaEval excluded).
Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg.
GPT-4o-Realtime Closed-source 74 51.6 71.7 69.7 66.7
Step-Audio-Chat Text-Driven 130B 43 56.1 56.7 36.7 48.1
MiniCPM-o 2.6 Text-Driven 8B 51 40.0 61.0 40.2 48.1
Llama-Omni Text-Driven 8B 39 22.9 45.3 10.7 29.5
VITA-1.5 Text-Driven 7B 20 28.1 46.7 23.3 29.5
GLM-4-Voice Interleaved 9B 51 32.0 50.0 36.4 42.3
GLM-4-Voice-Small Interleaved 1.5B 5.4 18.3 4.6 9.4
LUCY (S2) Parallel 7B 25.6 54.0 22.8 34.1
Moshi Parallel 7B 24 23.8 43.7 16.7 27.1
Mini-Omni Parallel 0.5B 25 12.8 22.0 6.9 16.7
OmniDRCA Parallel 8B 45 30.0 65.0 32.9 43.2
OmniDRCA-Small Parallel 2B 38 22.1 55.3 17.9 33.3
Table 3: Speech-to-Text ($S\rightarrow T$) Generation Performance on OpenAudioBench using benchmark-specific metrics (higher is better). Baseline results are from KimiTeam et al. (2025), except LUCY (S2) and Moshi from Gao et al. (2025) and MiniCPM-o 2.6 from our experiments. Reasoning QA is removed since our model is only trained on English data. Avg. is averaged over the four benchmarks (Gray: AlpacaEval excluded).
Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
Qwen2-Audio-Chat Cascaded 7B 57.2 45.2 69.7 40.3 53.1
MiniCPM-o 2.6 Text-Driven 8B 64.1 69.2 78.0 63.0 68.5
Step-Audio-chat Text-Driven 130B 56.5 73.0 72.3 56.8 64.7
Baichuan-Omni-1.5 Interleaved 7B 59.6 58.7 74.3 55.4 62.0
GLM-4-Voice Interleaved 9B 57.9 55.4 76.0 51.8 60.3
Moshi Parallel 7B 26.6 62.3 22.8 37.2
LUCY (S2) Parallel 7B 29.3 57.3 23.3 36.6
OmniDRCA Parallel 8B 58.0 51.7 79.7 47.7 59.7
Table 4: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following (Zeng et al., 2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text. Baseline scores are copied from (Zeng et al., 2025), except MiniCPM-o 2.6 results are from our experiments.
[…]
MiniCPM-o 2.6 Text-Driven 8B 4.18 13.31
Llama-Omni Text-Driven 8B 3.92 9.18
GLM-4-Voice Interleaved 9B 4.33 7.83
Moshi Parallel 7B 3.90 7.95
Mini-Omni Parallel 0.5B 3.17 25.28
OmniDRCA Parallel 8B 3.86 6.75
OmniDRCA-Small Parallel 2B 3.87 8.86
Table 5: Ablation study of Dual-Resolution Grouping Strategy Conducted on OmniDCRA-Small with Llama Questions Dataset.
Modal Group Size S2T T2M (T/S) T2T
S2T+T2T (Understanding) 1 55.33 61.33
5 62.33 69.67
T2M+T2T (Generation) 1 51.67 / 41.00 58.33
5 42.00 / 25.33 64.00
Table 6: Ablation study of Contrastive Cross-modal Alignment (Cont), SRM, and Multimodal Architecture Design (MD) on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
Model S2M (T/S) S2T T2M (T/S) T2T STC (T/S) SAC (T/S) SUC (T/S)
OmniDRCA-Small 64.33 / 55.33 64.67 69.67 / 58.67 69.67 67.33 / 59.67 65.67 / 57.00 67.67 / 57.00
w/o Cont 61.67 / 53.00 62.33 70.00 / 60.00 74.00 69.33 / 61.00 63.00 / 55.00 66.33 / 58.67
   w/o SRM 21.67 / 15.33 56.00 45.22 / 35.00 73.00 64.33 / 50.67 55.67 / 42.33 40.33 / 27.67
   w/o MD 58.00 / 49.00 58.00 69.33 / 55.00 68.33
4.1 Experimental Setup
Datasets. Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for SRM pre-training. 17.8K hours of English speech conversations data synthesized from text are adopted for OmniDRCA training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^2] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b). Following prior works (Yao et al., 2024; Li et al., 2025), we evaluate the performance on the widely used Spoken Question Answering benchmarks, including AlpacaEval (Li et al., 2023), Web Questions (Berant et al., 2013), Llama Questions (Nachmani et al., 2024), and TriviaQA (Joshi et al., 2017). To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, we adopt datasets from two established evaluation frameworks: UltraEval-Audio  [^3] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench  [^4] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation. As discussed in Huang et al. (2025); Chen et al. (2025), the results on TriviaQA are for reference only and marked with “*” due to the absence of an official test set. Moreover, we detect performance anomalies of OmniDRCA on the UltraEval-Audio TriviaQA dataset; therefore, we replace UltraEval-Audio TriviaQA dataset with OpenAudioBench TriviaQA dataset for a more meaningful evaluation  [^5] For reference, this setup yields a score of 42.6 for MiniCPM-o 2.6 while its score on the original UltraEval-Audio TriviaQA is 40.2..
Evaluation Metrics. Following the settings of UltraEval-Audio and OpenAudioBench, we utilize G-Eval (Liu et al., 2023) to evaluate model performance on AlpacaEval, and employ accuracy to assess model performance on other datasets. The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al. (2025).
Baselines. Representive baselines are selected to cover diverse modeling paradigms: the closed-source model GPT-4o-Realtime (OpenAI, 2024b), cascaded system Qwen2-Audio-Chat (Chu et al., 2024), Text-Driven models including Step-Audio-Chat (Huang et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni (Fang et al., 2024), and VITA-1.5 (Fu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025) and GLM-4-Voice (Zeng et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Mini-Omni (Xie and Wu, 2024a). This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
Implementation Details. SRM uses Qwen2.5-0.5B as its base LLM, trained following $T2M$ paradigm. OmniDRCA uses Qwen2.5-7B-Instruct as its base LLM, while OmniDRCA-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. The warmup rate is set to $1\%$ of the total training steps. For OmniDRCA, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for OmniDRCA-Small and SRM it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$. The AdamW method (Loshchilov and Hutter, 2019) is used for optimization. Experiments are run on 32x NVIDIA Tesla A100 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding. For OmniDRCA, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRM pretraining, and about $20$ hours on OmniDRCA post training.
4.2 Results and Discussions
Tables 2 and 3 compare OmniDRCA with the representative and competitive baselines across all major model categories, on speech-to-speech and speech-to-text generation performance. Table 2 shows that in prior works, best Text-Driven models substantially outperform top-tier Joint Speech-Text models, because text-driven systems do not allow speech tokens to interfere with the generation process, thereby minimally impacting the LLM’s generation paradigm. Moreover, prior interleaved models often outperform parallel models, though they exhibit higher latency (e.g., 13 text tokens for first speech in GLM-4-Voice) and lack temporal alignment, complicating multistream duplex modeling. As shown in Tables 2 and 3, OmniDRCA achieves SOTA performance under the parallel paradigm, achieving a substantial 24.9% relative gain over LUCY (S2) on the average score across three Speech-to-Speech QA benchmarks and 60.4% gain over Moshi across three Speech-to-Text QA benchmarks, excluding AlpacaEval in both cases. OmniDRCA also shows competitive performance against the interleaved GLM-4-Voice and Baichuan-Omni-1.5. However, a relatively large gap still remains when compared to text-driven systems. In future work, we plan to train the model on more extensive datasets to improve the performance.
The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (WER(%)), as shown in Table 4. Among models with comparable sizes, text-driven models tend to exhibit higher WER, while joint speech-text models achieve relatively lower WER. The model size also significantly impacts WER. OmniDRCA achieves the lowest WER, indicating strong alignment when generating both speech and text simultaneously. In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a). The gap in UTMOS between OmniDRCA and GLM-4-Voice suggests room for improvement in modeling prosody and preserving acoustic details.
We conduct extensive ablation study to analyze effects of critical algorithmic designs.
Dual-Resolution Grouping Strategy. Table 5 reveals the dual effects of our grouping strategy: while improving speech understanding (12.6% relative gain in S2T with group size 5) and text capabilities (13.6% relative gain in T2T), it degrades speech generation quality by 38.2% relative in T2M. On the other hand, as shown in Table 6, SRM demonstrates remarkable effectiveness in speech token generation, achieving 245.9% and 71.4% relative improvements in S2M and T2M tasks respectively. Notably, SRM also enhances speech comprehension metrics (S2T and STC-T) by 10-16% relative, suggesting that improved speech token reconstruction capability improves latent speech representations. The text-only performance (T2T) remains stable, confirming SRM’s modality-specific effectiveness without interfering with text processing pathways. Our dual-resolution architecture resolves this conflict by applying grouping only in the comprehension phase (5Hz) while using raw frame resolution (25Hz) with SRM for generation.
Contrastive Cross-modal Alignment. As shown in Table 6, the introduction of speech-text contrastive learning for sentence-level semantic alignment substantially enhances the model’s semantic understanding capabilities. While the final evaluation metric (S2S performance) shows improvement over the counterpart without contrastive loss, the enhancement exhibits modality-specific efficacy. For instance, S2T performance improves by 2.34% absolute (64.67 over 62.33), while T2T performance degrades by 4.33%, suggesting that while improving speech comprehension, the contrastive objective partially compromises the LLM’s native text understanding capacity. Our multimodal evaluation framework effectively reveals these nuanced impacts, guiding future work on mitigating text performance degradation through more balanced loss design.
Multimodal Architecture Design (MD). Our multimodal design achieves two key goals: (1) context-aware modality selection and (2) implicit data augmentation. As shown in Table 6, on the speech generation task, STC/SAC/SUC demonstrate substantial improvements over S2M. STC achieves 61.00, substantially surpassing the direct generation baseline of 55.33. This reveals the critical role of CoM while also indicating potential optimization opportunities for our S2M model. Without incorporating the three CoM variants (STC/SAC/SUC), the model fails to adopt this generation paradigm guided by system prompts. Final S2M performance shows that training without MD results in a 7.5% performance degradation, confirming the data augmentation effect of our approach.
Inspired by Wang et al. (2025), to enable full-duplex interaction, we propose a time-division multiplexing (TDM) input stream that allows OmniDRCA to accept user speech inputs during the assistant’s speech generation phase, thereby efficiently utilizing otherwise idle time slots. The resulting OmniDRCA-Duplex achieves full-duplex dialogue capabilities. The methodology and empirical validations of OmniDRCA-Duplex are detailed in Appendix.
5 Conclusions, Limitations, and Future Work
We introduce OmniDRCA, a novel parallel speech-text foundation model that integrates joint autoregressive modeling with dual-resolution speech representations and contrastive cross-modal alignment. Experiments show that OmniDRCA sets new SOTA speech-to-speech/text generation performance among parallel models and reaches competitive performance to interleaved models.
Limitations: (1) OmniDRCA’s speech generation quality remains constrained by suboptimal speech tokenization. We will explore other tokenization methods, such as CosyVoice’s speech tokenizer, to address this limitation. (2) We will extend exploration from current speech understanding and generation tasks to broader audio comprehension/synthesis tasks, and integrate visual modality.
Appendix A OmniDRCA-Duplex
As mentioned in Introduction, we explore the potential of OmniDRCA for full-duplex voice conversation. By leveraging time-division multiplexing, we introduce a novel spoken dialogue model, OmniDRCA-Duplex, which supports real-time responses to user interruptions with minimal performance degradation. In this section, we first introduce OmniDRCA-Duplex in Appendix A.1, then present the experimental results and discussions of its full-duplex conversation capabilities in Appendix A.2.
A.1 Methodology
As shown in Figure 2, compared with OmniDRCA, the overall architecture of OmniDRCA-Duplex remains largely unchanged, which is also composed of the multimodal LLM (MLLM) and the speech refinement model (SRM). The primary differences between OmniDRCA-Duplex and the original OmniDRCA model lie in two aspects. The first aspect is the organization of the input structure. OmniDRCA-Duplex organizes user speech and assistant speech alternately along the temporal axis as input, referred to as the Time-Division Multiplexing Input Stream, while OmniDRCA adopts turn-based speech as the same as text-based LLM. For the second aspect, to enhance the model’s sensitivity to user interruptions, an additional assistant state prediction head is introduced in OmniDRCA-Duplex to determine whether muting should be activated. The following paragraphs provide detailed descriptions of these two designs.
Figure 2: The architecture of OmniDRCA-Duplex.
Time-Division Multiplexing Input Stream.
Full-duplex spoken dialogue models allow simultaneous receptions of inputs from both user and assistant. Inspired by Wang et al. (2025), OmniDRCA-Duplex organizes inputs into a double-channel format, specifically, an alternating sequence of user speech and assistant speech along the temporal axis. Besides, same position id is used in one time slot, which paired user and assistant speech representations together. The model enforces attention masking between concurrent speech representations in individual time slots since they appear simultaneously. This structured arrangement, termed time-division multiplexing (TDM) input stream, is visually illustrated in Figure 2 TDM-Duplex. This approach allows the model to accept user speech when the assistant is generating speech, effectively utilizing the time slots that would otherwise be idle. The TDM input stream is designed to handle both user and assistant speech inputs, ensuring a seamless interaction experience.
Awareness of Interruptions.
As our early experiments show, the model may not be aware of user interruptions during the generation of assistant speech, since the model prefers to generate speech based on the content in the time slots from its own role (that is, Assistant Speech Token Stream). This can lead to a lack of responsiveness to user interruptions and a less engaging interaction experience. To address this issue, we implement a mechanism that enhances the model’s awareness to user interruptions during assistant speech generation. We achieve this goal by adding an additional head to predict the assistant’s speech generation state. The model is trained to predict whether the assistant is currently generating speech or not, based on the input speech tokens. This additional head provides the model with contextual information about the assistant’s speech generation state, enabling it to respond more effectively to user interruptions.
A.2 Experiments
Experimental Setup.
The original speech datasets using for OmniDRCA post-training are edited to simulate full-duplex conversation scenarios. Two types of datasets are constructed: 1) Non-interruption Dataset: Dialogue turns are structured into dual-channel audio. To better simulate human interactions, a response time of $0\sim 400$ ms is randomly inserted between dialogue turns. This dataset enables the model to learn when to initiate and terminate speech. 2) Interruption Multi-Turn Dialogue Dataset: The assistant’s responses in dialogue turns are randomly truncated, followed by inserting a new user input , before constructing dual-channel datasets. Through this data simulation approach, the model learns to handle incomplete interrupted semantics. In total, about 4K hours non-interruption dataset and 4K hours interruption multi-turn dialogue dataset are adopted for OmniDRCA-Duplex training. UltraEval-Audio Benchmark is used for evaluating the semantic ability, while multi-turn Llama Questions randomly pairing two of questions together is used for evaluating the interruption ablitily. Specific, providing these two rounds Llama Questions examples, if the model can mute itself while the user is speaking and respond after the user finishes, it is considered a successful Turn-taking. The testing evaluates the success rate of Turn-taking. 666 Here, Turn-taking@k Metrics is not conducted unlike other duplex speech models (Zhang et al., 2024a). This approach stems from intentional dataset construction aimed at endowing the model with human-like response timing and enabling scenarios where semantic information coexists simultaneously in both channels. The design specifically emphasizes making the model’s turn-taking response time more flexible rather than prioritizing the speed of turn-taking as the objective for the duplex model.
Table 7: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks Comparing the original OmniDRCA model and its Duplex Model.
Model Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
Moshi 7B 24 23.8 43.7 16.7 27.1
OmniDRCA-Duplex-Small 2B 33 16.1 46.6 13.6 27.3
OmniDRCA-Small 2B 38 22.1 55.3 17.9 33.3
Results.
As demonstrated in Table 7, evaluation results on the UltraEval-Audio Spoken Question Answering Benchmark indicate that OmniDRCA-Duplex-Small attains an average score of 27.3. This represents a 18.2% relative performance degradation compared to the baseline OmniDRCA-Small model’s average score of 33.3. Notably, the experimental results show that OmniDRCA-Small-Duplex (2B model size) has a slight performance advantage over the larger 7B full-duplex model Moshi (27.3 versus 27.1). For the model’s interruption performance, by setting the multi-turn Llama Questions with intervals of $t=3s$ between turns, OmniDRCA-Small-Duplex achieves 98.64% success rate of Turn-taking, meaning the model can keep silent when user is talking and consistently respond to user-interrupted utterances in the second round.
Table 8: The Impact of SRM Pretraining Data on Llama Questions Performance.
Model S2M (T/S) S2T T2M (T/S) T2T
OmniDRCA-Small (Less) 52.67 / 44.33 60.33 67.67 / 53.00 75.00
   w/o. SRM-PT 38.33 / 30.33 56.00 59.33 / 46.33 73.33
Figure 3: Performance Scaling of OmniDRCA-Small (w/o. Cont) on LLaMA Question Benchmark.
Appendix B Dataset Scaling
Pre-Training
To examine the impact of pretraining data on model performance, the SRM pretraining step is eliminated, and the OmniDRCA-Small model is retrained. For computational efficiency, fewer post-training steps are adopted, called OmniDRCA-Small (Less), leading to the inferior performance compared to the results of OmniDRCA-Small reported in the main text. As shown in Table 8, SRM pretraining has the most significant impact on the performance of S2M and T2M, while its effect on S2T and T2T is relatively minor. This is reasonable because SRM is primarily responsible for refining speech token generation tasks, thus exerting a smaller influence on purely textual generation. This underscores the importance of SRM pretraining.
Post-Training
To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on OmniDRCA-Small (w/o. Cont). Each S2S instance can be augmented into 7 distinct multimodal interaction patterns. As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements. The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
Appendix C Prompt Templates
System prompts for multimodal interaction patterns.
As shown in Table 9, there are five types of system prompts categorized based on the model’s output partitioning. During inference, specifying these system prompts enables the generation of desired output results.
Table 9: System Prompts for Multimodal Interaction.
Pattern Abbr. System Prompts
S2M & T2M You are a helpful assistant and asked to generate both text and speech tokens at the same time.
S2T & T2T You are a helpful assistant and asked to generate text tokens.
STC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, think of an appropriate text response, and then convert the response back to both text and speech tokens at the same time.
SAC You are a helpful assistant. Let’s think step by step. Think of an appropriate text response, and then convert the response back to both text and speech tokens at the same time.
SUC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, and then think of both appropriate text and speech responses at the same time.
Evaluation Templates.
To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, datasets from two established evaluation frameworks were adopted: UltraEval-Audio [^7] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench [^8] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation. For UltraEval-Audio, string-matching algorithms were applied to Llama Questions, Web Questions, and TriviaQA, while AlpacaEval utilized GPT-4o-mini with the Alpaca scoring template. In OpenAudioBench, GPT-4o-2024-08-06 was employed for scoring with different templates shown in Figures, all in strict compliance with the tools’ prescribed methodologies.
AlpacaEval Evaluation Template [Instruction] Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: "[[rating]]", for example: "Rating: [[5]]". [Question] (Instruction Transcription) [The Start of Assistant’s Answer] (Response Transcription) [The End of Assistant’s Answer]
TriviaQA and Web Questions Evaluation Template Your will be given a question, the reference answers to that question, and an answer to be judged. Your tasks is to judge whether the answer to be judged is correct, given the question and reference answers. An answer considered correct expresses or contains the same meaning as at least **one of** the reference answers. The format and the tone of the response does not matter. You should respond in JSON format. First provide a one-sentence concise analysis for the judgement in field ‘analysis’, then your judgment in field ‘judgment’. For example, “‘json "analysis": "<a one-sentence concise analysis for the judgement>", "judgment": < your final judgment, "correct" or "incorrect"> ”’ # Question (instruction) # Reference Answer (targets) # Answer To Be Judged (answer_to_be_judged)
Llama Questions Evaluation Template ## Background You are a professional QA evaluation expert. You need to assess whether the model’s answer is correct based on the standard answer. ## Scoring Criteria Correct: The answer matches or is equivalent to the standard answer Incorrect: The answer is wrong or irrelevant to the question ## Evaluation Guidelines 1. The expression of answers can be flexible, not requiring exact matches. For example: - Numbers can be expressed in either Arabic numerals or words - Proper nouns can be in either English or Chinese - Differences in punctuation can be ignored 2. Focus on whether the core meaning of the answer is correct ## Output Format Provide the reasoning for your score, then generate the result in "[]" format and make sure it contains "the score is [Correct]" or "the score is [Incorrect]", for example: “‘ The answer is correct and equivalent to the standard answer, the score is [Correct] ”’ or “‘ The answer is incorrect and does not match the standard answer, the score is [Incorrect] ”’ ## Question: (instruction) ## Standard Answer: (targets) ## Model’s Answer: (answer_to_be_judged)