### v2

# DrVoice: Parallel Speech-Text Voice Conversation Model via Dual-Resolution Speech Representations

Abstract
Recent studies on end-to-end speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens. Existing approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis. (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation. This paper presents DrVoice, a parallel speech-text voice conversation model based on joint autoregressive modeling, featuring dual-resolution speech representations. Whereas current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz. Experimental results on Spoken Question Answering benchmarks demonstrate that DrVoice establishes new state-of-the-art (SOTA) performance among similar size speech foundation models with relative small amount of data.
1 Introduction
Recent advancements in Large Language Models (LLMs) (OpenAI, 2023; Yang et al., 2024) have demonstrated remarkable performance across various NLP tasks, sparking growing interest in extending their capabilities to multimodal domains encompassing vision and audio. Developments in spoken dialogue systems are critical to human-computer interaction, as natural human communication inherently relies on verbal exchanges. The release of GPT-4o (OpenAI, 2024b) is a significant milestone in this field, showing the potential of LLM-based spoken dialogue systems for seamless and natural interactions with users.
LLM-based spoken dialogue systems can be generally categorized into cascaded systems and end-to-end (E2E) systems, with the distinction lying in whether the backbone LLM can directly comprehend speech representations and generate speech outputs. Early cascaded systems integrate separately trained Automatic Speech Recognition (ASR), LLM, and Text-to-Speech (TTS) modules. Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., tone, emotion), and significant latency. Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (Chu et al., 2023, 2024). E2E systems have emerged to further eliminate ASR and TTS modules, and establish direct connections between speech representations and LLMs. However, training LLMs to generate highly intelligent speech outputs remains challenging. E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
Recent research efforts on E2E models to address the above challenges have focused on two primary directions Chen et al. (2024): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models. Text-Driven Speech Models refer to systems in which LLMs process speech representations as input, produce textual responses, and utilize the LLM’s representations as input to a speech decoder for speech generation. In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and generating both text tokens and speech tokens simultaneously. The key distinction lies in whether the generation of speech tokens can influence the subsequent generation of text tokens within the LLM. This feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models. Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024; KimiTeam et al., 2025). Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM. While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations. The architecture of Text-Driven Speech Models creates a unidirectional information flow where text generation is completed before speech synthesis begins. This prevents the LLM from conditioning its textual output on the generated acoustic tokens, thus limiting its ability to control fine-grained paralinguistic attributes like emotion and prosody.
Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge. Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials Chen et al. (2024); hence, in this work, we focus on enhancing joint speech-text modeling. Recently, Kimi-Audio KimiTeam et al. (2025) has set a new state-of-the-art (SOTA) in joint speech-text modeling. However, this approach faces notable challenges. It not only requires extensive data for training but also incurs significant computational costs due to its 12.5Hz audio representation. Furthermore, this high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR), which compresses the input down to 5Hz. DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens. For speech comprehension, we introduce a grouping mechanism that maps 25Hz discrete audio tokens to 5Hz speech representations, which better match the pace of text. During generation, combined speech-text embeddings form the assistant’s autoregressive input. The hidden states captured from shared LLM layers are then passed in parallel to a text head for text token prediction, and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens. Due to input-output resolution discrepancies, each LLM inference triggers $k$ parallel SRH executions, where $k$ is the grouping factor. To further enhance the model’s reasoning and the coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy. This approach prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation. This intermediate reasoning step serves to improve modality alignment and reduce errors. For practical applications, we design system prompts to control the output modes, enabling text-only output, direct parallel speech-text output, or the CoM-enhanced parallel output. Experimental results demonstrate the efficacy of CoM in significantly improving performance. For reproducibility, we will release all source code and model checkpoints.
Our contributions are three-fold: 1) We propose DrVoice, a novel parallel speech-text conversation model featuring Dual-Resolution Speech Representations (DRSR). This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs and better preserving the LLM’s semantic processing capabilities. 2) We introduce a CoM-Mixing training strategy which acts as a curriculum, using the structured reasoning of CoM to scaffold directly speech generation task. 3) Through comprehensive experiments, we demonstrate that DrVoice achieves new state-of-the-art (SOTA) performance on multiple spoken question answering benchmarks, validating the effectiveness of our proposed architecture and strategies.
2 Related Work
Speech Tokenization.
Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations. Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation. In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as in CosyVoice (Du et al., 2024a, b), and prioritized for linguistic content (Hsu et al., 2021). While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023). In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to their competitive semantic capabilities and compatibility with a more lightweight, diffusion-free vocoder for direct speech synthesis.
E2E Speech Foundation Models.
Existing E2E speech foundation models can be categorized into Text-Driven Speech Models and Joint Speech-Text Models. Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech. Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production. Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
Joint Speech-Text Models explore two different architectures, Interleaved and Parallel. Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024; KimiTeam et al., 2025) conduct parallel decoding. The interleaved model VITA-Audio Long et al. (2025) achieves efficient audio generation by introducing multiple cross-modal token prediction (MCTP) modules. It demonstrates superior performance across multiple benchmarks while maintaining high-quality speech synthesis. The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations. To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper Radford et al. (2023) features, preserving both semantic and acoustic information. In order to further enhance the efficiency of both training and inference and alleviate the issue of granularity misalignment between the text and speech modalities, our approach leverages DRSR to reduce the LLM input frame rate to 5Hz, without sacrificing performance.
3 Methodology
Figure 1 illustrates the architecture of DrVoice. The system consists of three main components: (1) Speech Tokenizer and Speech Encoder are designed to process the wave of speech into hidden representation, (2) a Multimodal Large Language Model (MLLM) consists of shared LLM layer, text head, and speech refined head (SRH) for token generation, and (3) Speech Detokenizer to generate wave from speech tokens. The system operates through multimodal input encoding and coordinated speech-text output generation. During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce parallel speech-text responses through SRH and text head. To effectively train this system, we introduce the CoM-Mixing Training strategy.
3.1 Speech Tokenization and Detokenization
Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ HuBERT (Hsu et al., 2021) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$, where $T$ is the speech token sequence length. For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency. Inspired by Kimi-Audio KimiTeam et al. (2025) and aiming for improved audio understanding, we utilize the Whisper-Large-v3 Speech Encoder to extract continuous audio representations. Subsequently, an Adapter is introduced to downsample the temporal resolution of these representations and align their hidden dimension with that of the LLM. Finally, the processed continuous representations are stacked with discrete ones and fed into the LLM.
Figure 1: Overview of DrVoice. User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text and speech token prediction. The generated speech tokens are then converted to speech waveform by the speech detokenizer. Note that SRH through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor.
Table 1: Multimodal Interaction Patterns.
Pattern Name Abbr. Modality Flow
Speech-to-Multimodal S2M Speech $\rightarrow$ Joint speech-text response
Speech-to-Text S2T Speech $\rightarrow$ Text-only response
Text-to-Multimodal T2M Text $\rightarrow$ Joint speech-text response
Text-to-Text T2T Text $\rightarrow$ Text-only response
Speech-Text Chain STC Speech $\rightarrow$ Text transcription $\rightarrow$ Text response $\rightarrow$ Multimodal response
Speech-Assisted Chain SAC Speech $\rightarrow$ Text response (agent perspective) $\rightarrow$ Multimodal response
Speech-User Chain SUC Speech $\rightarrow$ Text transcription (user perspective) $\rightarrow$ Multimodal response
3.2 Multimodal Large Language Model (MLLM)
Built upon text-LLMs, the MLLM learns joint speech-text modeling to process speech or text inputs while producing parallel speech and text outputs.
Parallel Joint Speech-Text Model.
Leveraging the autoregressive generation capability of LLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the shared LLM layer at each timestep. Their embeddings are added as model input, forming a parallel speech-text architecture. Formally, the combined input embedding $c_{t}$ at timestep $t$ is computed as:
$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. Since the lengths of speech tokens and text tokens are typically mismatched, the shorter sequence is padded with special token <|SIL|>to align them for each utterance. The autoregressive generation process is as follows:
$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
where $x$ is the input sequence and $y_{t}=(s_{t},t_{t})$ denotes the joint speech-text output at timestep $t$. This unified formulation enables seamless integration of speech and text generation within a single autoregressive framework. To preserve the intrinsic linguistic understanding and generation capabilities of pretrained text-LLMs while enabling cross-modal interactions, three key methodological components are introduced for intermodal coordination.
Inner Text Streaming.
Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold. We focus on performing modality alignment exclusively at the assistant end. This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
Speech Token Grouping.
A crucial parameter for discrete tokens is sampling rate, which determines the input/output sequence length. GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz. Hence, our work adopts 25Hz sampling rate. To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024), a grouping mechanism is designed:
$\mathbf{g}_{i}=\text{Linear}\left(\mathop{\|}_{j=ik}^{(i+1)k-1}\mathbf{s}_{j}\right)\in\mathbb{R}^{d_{\text{text}}}$ (3)
where $\mathbf{s}_{j}$ denotes speech tokens, $\|$ represents feature concatenation, and $k$ is the grouping factor determined by the ratio between speech token rates and text token rates. This design compresses the speech token length from $T$ to $T/k$ and the grouped speech representations align better with text. Different from Chen et al. (2024) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
Speech Refined Head.
SRH is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens. While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details. To address this limitation, an ungrouping process is performed as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[SLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (4)
and time splitting
$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}],$ (5)
where $\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k}$, the resulting $\mathbf{H}$ serves as conditional input for SRH that autoregressively generates speech tokens. Following our practice on SLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated. Speech token prediction is trained to maximize the conditional probability:
$\mathcal{L}_{\text{SRH}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}_{<i}),$ (6)
where $s_{i}$ represents the $i$-th speech token. By optimizing this objective, SRH learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from SLLM. This design enables SRH to effectively leverage the semantic and acoustic information encoded in the hidden representations of SLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
The end-to-end training objective integrates the two losses through multi-task learning paradigm:
$\mathcal{L}_{\text{MLLM}}=\lambda\mathcal{L}_{\text{TH}}+\mu\mathcal{L}_{\text{SRH}},$ (7)
$\mathcal{L}_{\text{TH}}=-\sum_{i=1}^{T}\log P(t_{i}|c_{<i},\mathbf{g}),$ (8)
where $\mathcal{L}_{\text{TH}}$ is the autoregressive loss over text head, $\lambda$ and $\mu$ are hyperparameters.
3.3 CoM-Mixing Training strategy
The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in the Appendix A. During inference, specifying these system prompts enables the generation of desired output results. For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
4 Experiments
Table 2: Speech-to-Text (S $\rightarrow$ T) and Speech-to-Speech (S $\rightarrow$ S) generation performance on spokenQA benchmarks. The results for the S2T baselines are taken from the previous SOTA model Kimi-Audio’s paper (Xu et al., 2025), and the S2S results are taken from the previous SOTA model VITA-Audio’s paper (Long et al., 2025). Otherwise, results are taken from their original papers. The model size refers only to the parameter count of the LLM backbone, excluding the audio encoder, audio tokenzier, and audio detokenzier module. Frame Rate (In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone. The average number of text tokens corresponding to one second of speech is denoted by $\tau$. Metrics are benchmark-specific (higher is better). The Avg. score averages two benchmarks. The best results in each group are highlighted in bold, while the second-best results are underlined.
Model Size Frame Rate (In/Out) Llama Q. Web Q. Avg.
$\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$
[…]
DrVoice-Small (Our) 2B 5/5 73.0 56.0 42.1 31.1 57.5 43.5
Moshi (2024) 7B 12.5/12.5 62.3 21.0 26.6 9.2 44.5 15.1
[…]
LUCY (S2) (2025) 7B 12.5/48 57.3 51.0 29.3 18.2 43.3 34.6
[…]
Kimi-Audio (2025) 8B 12.5/12.5 79.3 - 70.0 - 74.7 -
[…]
Stream-Omni (2025) 7B 25/25 76.3 65.0 44.2 27.5 60.3 46.3
[…]
DrVoice (Our) 8B 5/5 85.0 71.0 65.2 40.1 75.1 55.5
Table 3: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following Zeng et al. (2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text. Baseline scores are copied from Zeng et al. (2025). FR(In/Out) indicates the input speech frame rate and the output (speech + text) frame rate for the LLM backbone, while $\tau$ represents the average text tokens per second of speech. The best results in each group are bolded.
[…]
Mini-Omni (2024a) 0.5B 50/48 3.17 25.28
DrVoice-Small 2B 5/5 3.87 8.86
[…]
Moshi (2025) 7B 12.5/12.5 3.90 7.95
[…]
DrVoice 8B 5/5 3.86 6.75
Table 4: Impact of the Grouping Factor on Llama Q. S2T and S2M performance. Experiments were conducted on DrVoice-Small w/o. CSE, using data subsets for faster exploration. The S2T model was trained exclusively on mixed S2T and T2T data, while the S2M model was trained on mixed S2M and T2T data.
Grouping Factor S2T S2M (T/S)
1 55.67 4.00 / 2.67
3 64.67 15.67 / 5.00
5 63.33 37.67 / 28.00
7 62.67 36.00 / 16.67
Table 5: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing Training on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
Model S2M (T/S) S2T T2M (T/S) T2T STC (T/S) SAC (T/S) SUC (T/S)
DrVoice-Small 68.67 / 56.00 72.33 72.33 / 56.00 75.33 75.67 / 68.33 71.67 / 62.67 73.33 / 62.00
w/o. CSE 61.67 / 53.00 62.33 70.00 / 60.00 74.00 69.33 / 61.00 63.00 / 55.00 66.33 / 58.67
   w/o. SRH-Pretraining 38.33 / 30.33 56.00 59.33 / 46.33 73.33 67.33 / 57.67 54.00 / 42.33 54.33 / 42.67
    w/o. SRH 21.67 / 15.33 56.00 45.22 / 35.00 73.00 64.33 / 50.67 55.67 / 42.33 40.33 / 27.67
   w/o. CoM-Mixing 58.00 / 49.00 58.00 69.33 / 55.00 68.33 –/– –/– –/–
Figure 2: Computational Resources across different Grouping Factor.
Figure 3: Performance Scaling of DrVoice-Small (w/o. Continuous Speech Encoder) on LLaMA Question Benchmark.
4.1 Experimental Setup
Datasets.
Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for pre-training speech refined head. 17.8K hours of English speech conversations data synthesized from text are adopted for DrVoice training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^1] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b). Following the common practice in prior works (Long et al., 2025; KimiTeam et al., 2025; Yao et al., 2024; Zhang et al., 2025), the Speech-to-Speech ($S\rightarrow S$) and Speech-To-Text ($S\rightarrow T$) performance is evaluated on widely used Spoken Question Answering benchmarks, including Web Questions (Berant et al., 2013) and Llama Questions (Nachmani et al., 2024).
Evaluation Metrics.
Accuracy is emplyed to assess model performance on datasets. The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al. (2025).
Baselines.
Representive baselines are selected to cover diverse modeling paradigms: the cascaded system Audio Flamingo 3 (Zhang et al., 2025), Text-Driven models including Freeze-Omni (Défossez et al., 2024), MinMo (Chen et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni2 (Fang et al., 2025), and Qwen2.5-Omni (Xu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025), GLM-4-Voice (Zeng et al., 2025), Stream-Omni (Zhang et al., 2025), and VITA-Audio (Long et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Kimi-Audio (Xu et al., 2025). Noting that Kimi-Audio achieves state-of-the-art (SOTA) performance on Speech-to-Text Question Answering, while VITA-Audio sets the SOTA on Speech-to-Speech Question Answering. This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
Implementation Details.
Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT). DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. The warmup rate is set to $1\%$ of the total training steps. For DrVoice, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for DrVoice-Small and SRH-PT it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$. The AdamW method (Loshchilov and Hutter, 2019) is used for optimization. Experiments are run on 32x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding. For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post training.
4.2 Results and Discussions
[…]
The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (ASR-WER(%)), as shown in Table 3. Among models with comparable sizes, DrVoice achieves low ASR-WER, indicating strong alignment when generating both speech and text simultaneously. In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a). The gap in UTMOS between DrVoice and GLM-4-Voice suggests room for improvement in speech naturalness and intelligibility.
We conduct extensive ablation study to analyze effects of critical algorithmic designs.
Continuous Speech Encoder.
As shown in Table 5, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs. Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation. Specifically, the S2T score decreases from 72.33 to 62.33 (13.8% relatively), and the S2M (T) score falls from 68.67 to 61.67 (10.2% relatively). In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00. This confirms the CSE’s effectiveness in processing and representing speech information.
Dual-Resolution.
Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation.
Grouping Factor. Table 4 and Figure 2 demonstrate that our grouping strategy significantly enhances both performance and computational efficiency. Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M). For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain). The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5. Furthermore, as shown in Figure 2, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, suggesting the efficiency of grouping machinism.
Speech Refinement. As shown in Table 5, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks. By comparing the model with SRH (w/o. SRH-Pretraining) to one without (w/o. SRH), we observe that adding SRH achieves a 76.9% relative improvement in S2M (T) (from 21.67 to 38.33) and a 31.2% relative improvement in T2M (T) (from 45.22 to 59.33). The text-only performance (T2T) remains stable, confirming that SRH enhances speech generation without interfering with text processing pathways. Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
To examine the impact of pretraining, we remove the SRH pretraining step and retrain the model. As shown in Table 5, removing pretraining (comparing w/o. CSE with w/o. SRH-Pretraining) has the most significant impact on speech generation, causing S2M (T) performance to drop by 37.8% and T2M (T) by 15.2%. The effect on S2T is smaller (10.2% drop), and it is negligible for T2T. This underscores the critical importance of SRH pretraining for refining speech token generation.
CoM-Mixing Training Strategy.
As shown in Table 5, where tasks guided by contextual system prompts (STC/SAC/SUC) demonstrate substantial improvements over direct S2M generation. For example, STC (T) achieves a score of 75.67, significantly surpassing the S2M (T) baseline of 68.67. This shows the model successfully learns to adopt the generation paradigm guided by system prompts. The data augmentation effect is confirmed by ablating the CoM-Mixing strategy entirely. Training without it (w/o. CoM-Mixing Training) results in a 15.5% relative performance degradation in S2M (T) (from 68.67 to 58.00), confirming the value of our mixed-modality training approach.
Data Scaling
To investigate the impact of data scaling, we conduct experiments by progressively expanding the S2S training data on DrVoice-Small (w/o. Continous Speech Encoder). Each S2S instance can be augmented into 7 distinct multimodal interaction patterns. As illustrated in Figure 3, expanding the dataset from 3.6K to 17.8K samples yields consistent improvements. The performance of T2M(S) exhibits nearly linear growth, while other patterns, despite showing slower growth rates, still demonstrate upward trends in the figure, suggesting potential for further performance enhancement through additional data.
5 Conclusions and Future Work
We introduce DrVoice, a novel parallel speech-text voice conversation model that leverages joint autoregressive modeling with dual-resolution speech representations. Experimental results demonstrate that DrVoice sets a new state-of-the-art (SOTA) benchmark in both speech-to-speech and speech-to-text generation tasks against current baselines. Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency without sacrificing performance.
Our future work will address three key areas: (1) Enhancing Speech Generation Quality: To overcome the limitations of the current suboptimal tokenization and further improve speech fidelity, we plan to explore advanced speech tokenizers, such as the one utilized in CosyVoice. This will enable the model to capture finer acoustic details and generate more natural-sounding audio. (2) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations. Inspired by recent advancements like Parrot (Wang et al., 2025), we will investigate the use of a time-division multiplexing (TDM) input stream. This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue. (3) Expanding to General Audio and Multimodality: Finally, we aim to extend our research beyond speech-centric tasks to broader audio comprehension and synthesis, including the recognition and generation of music and environmental sounds. The subsequent integration of the visual modality will be a critical step toward developing a more comprehensive, multimodal conversational AI.
Appendix A Prompt Templates
As shown in Table 6, there are five types of system prompts categorized based on the model’s output partitioning. During inference, specifying these system prompts enables the generation of desired output results.
Pattern Abbr. System Prompts
S2M & T2M You are a helpful assistant and asked to generate both text and speech tokens at the same time.
S2T & T2T You are a helpful assistant and asked to generate text tokens.
STC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, think of an appropriate text response, and then convert the response back to both text and speech tokens at the same time.
SAC You are a helpful assistant. Let’s think step by step. Think of an appropriate text response, and then convert the response back to both text and speech tokens at the same time.
SUC You are a helpful assistant. Let’s think step by step. Convert speech to text if the query is speech, and then think of both appropriate text and speech responses at the same time.
Table 6: System Prompts for Multimodal Interaction.